OpenStudy (anonymous):

Find a particular solution of the nonhomogeneous equation and give the general solution of the equation. y''-(1/x)y'+(1/x^2)y=2/x y1(x) = x and y2(x) = xlnx

OpenStudy (anonymous):

For a solution, i got z(x) = 2lnx + 2lnx(xlnx) but that seems like it would be wrong. So I doubt my general solution is correct either. Can someone please check this?

OpenStudy (experimentx):

this looks like Euler-Cauchy equation

OpenStudy (anonymous):

it doesnt really mention that in the book... it just kind of gives guidelines how to find the solutions. they use the wronskian and the equation z(x) = u(x) y1+v(x) y2 to find the solution and general solution.

OpenStudy (experimentx):

try substituting z=ln(x) ... you will have second order ODE with e^z in terms of y and z ... solving it quite easy ... just substitute back x from z

OpenStudy (anonymous):

you mean for my solution?

OpenStudy (experimentx):

no i haven't check it ... what method did you use to deduct it?

OpenStudy (experimentx):

According to Mathematica, the solution is \( x C[1]+x C[2] \text{Log}[x]+x \text{Log}[x]^2 \)

OpenStudy (anonymous):

well i calculated the wronskian first using my y1 and y2 and got W(x) = x then I calculated u(x) = integral (-y2)f(x))/Wx then I calculated v(x) = inegral (y1/f(x)/Wx then plugged into z(x) = u(x) y1 + v(x) y2

OpenStudy (experimentx):

I mean what method did you use to deduct y1(x) = x and y2(x) = xlnx ??

OpenStudy (anonymous):

oh, i didnt, it was given.

OpenStudy (anonymous):

we are just told to verify they are solutions

OpenStudy (experimentx):

the method it's asking is called variation of parameters http://en.wikipedia.org/wiki/Variation_of_parameters

OpenStudy (anonymous):

yes... that is it.

OpenStudy (experimentx):

in this picture, W is your wonskian, u1 and u2 are your given homegeneous equation's solution ... and put f(x)=1/x and evaluate those integrals.

OpenStudy (experimentx):

woops!! wrong picture http://upload.wikimedia.org/math/1/5/7/157873c4bc16e492763bda5edd135911.png

OpenStudy (experimentx):

and add them ... A(x) and B(x) ... and don't forget to add integration of constants.

OpenStudy (anonymous):

okay now I got z(x) = x(ln x)^2

OpenStudy (experimentx):

all right ... that's correct. according to W|A

OpenStudy (anonymous):

ohhh, well what is W|A, or whatever that was that you used?

OpenStudy (anonymous):

ohhh thats pretty cool... i have never tried that site... well great way to check my answers for sure.. thank you a bunch

OpenStudy (experimentx):

yw